A water pump lifts water from a level 10 m below the ground. Water is pumped at a rate of 30 kg/minute with negligible velocity. Calculate the minimum horsepower the engine should have to do this.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. h = 10 m flow rate = (m/t) = 30 kg/min = 0.5 kg/sec power P = mgh/t = (0.5) * 9.8 * 10 = 49 W So, horse power (h.p) p/746 = 49/746 = 6.6 * 10^-2hp